OpenStudy (anonymous):

A sock drawer contains 18 blue socks and 10 black socks. If two socks are randomly drawn from the drawer without replacement, what is the probability that two blue socks are drawn?

OpenStudy (campbell_st):

well on the 1st draw you have P(Blue) = 18/28 2nd draw if the 1st draw was blue P(Blue) = 17/27 so P(Blue, Blue) = 18/28 x 17/27

OpenStudy (anonymous):

ok Thanks so the it will be 153/378

OpenStudy (campbell_st):

makes sense to me